At NTP calculate the number of atom in 5.6 L diatomic gas

Grade:11

1 Answers

harsha
6 years ago
1molecule of diatomic gas contain 2 atomsAt NTP, Moles (n)=V(l)/22.4n= 15.6/22.4 =0.70.7=no. Of atoms/ Avogadro no0.7=2*N/6.022*10^23. N=2.11*10^23